# Small Gradient Norm Regret for Online Convex Optimization

### v2

@@ Appendix B Proof of results in Section 3
This completes the proof by doing a case analysis on the value of $\eta=\min\{\sqrt{\tfrac{{D(D+2\mathcal{P}_{T}})}{{2\hat{G}_{T}}}},\frac{1}{2L}\}$.
B.2 Dynamic regret and proof of results in Section 3.1
@@
which can be further bounded by taking any $\bar{x}\in\mathcal{X}$ and $\max_{t\in[T]}\max_{x\in\mathcal{X}}\|\nabla\ell_{t}(x)\|\leq\max_{t\in[T]}\|\nabla\ell_{t}(\bar{x})\|+LD$. This constant $M$ will appear in a $\log\log$ term in the final result.
@@ B.2.1 Algorithm design
input initial vector $p^{1}=\tfrac{1}{N}\textbf{1}_{N}$ and $\eta_{i}=\eta_{\min}2^{i-1},i\in[N]$, expotentiated gradient learning rate $\delta>0$, initial point $x^{1}_{i}=x^{1},i\in[N]$.
@@ B.2.3 Algorithm analysis
$\hat{\rho}(\{\hat{x}_{t}\}_{t=1}^{T})\leq 8\sqrt{(3+\log N){D^{2}}+2\mathcal{P}_{T}D}+8(3+\log N)LD^{2}+8LD\mathcal{P}_{T}=\mathcal{O}(\sqrt{(1+\mathcal{P}_{T}+\hat{G}_{T})(1+\mathcal{P}_{T})}\log\log T)$
@@ Appendix C Proof of results in Section 4
$T[f(\bar{x}_{T})-f(x^{\star})]\leq{}$ $\textstyle\mathbb{E}\big[\sum_{t=1}^{T}[f(x^{t})-f(x^{\star})]\big]$
@@
C.2 Proof of Theorem 4.1
@@
since $\eta^{2}-\tfrac{\eta}{L}\leq 0$ for $\eta\leq\tfrac{1}{L}$. Next we bound $\sum_{j\in\{k:k\geq 2,\eta_{k}\geq\tfrac{1}{L}\}}(\eta^{2}_{j}-\tfrac{\eta_{j}}{L})\|\nabla\ell_{j}(x^{j})\|^{2}$ by taking $a_{0}=\|\nabla\ell_{1}(x^{1})\|^{2}$, $a_{j}=\|\nabla\ell_{j+1}(x^{j+1})\|^{2}$ and $g(u)=\tfrac{\alpha^{2}}{u}-\frac{\alpha}{L\sqrt{u}}$ in Lemma˜A.1:
@@
C.3 Proof of Proposition 4.1
@@
$\textstyle\sum_{t=1}^{T}\ell^{1/\gamma}_{t}(x)-\inf\ell^{1/\gamma}_{t}={}$ $\textstyle\sum_{t=1}^{T}\ell^{1/\gamma}_{t}(x)-\min_{x\in\mathcal{X}}\ell_{t}(x)$
$\leq{}$ $\textstyle\sum_{t=1}^{T}\ell^{1/\gamma}_{t}(x)-\min_{x\in\mathcal{X}}\ell_{t}(x)=L^{\mathcal{X}}_{T}(x),$
where the equality uses the fact that $\inf\ell^{1/\gamma}_{t}=\min_{x\in\mathcal{X}}\ell_{t}(x)$ and the last inequality uses $\ell(x)\leq\ell^{1/\gamma}(x)$. This completes the proof.

### v3

@@ Appendix B Proof of results in Section 3
This completes the proof by doing a case analysis on the value of $\eta=\min~\{\sqrt{\tfrac{{D(D+2\mathcal{P}_{T}})}{{2\hat{G}_{T}}}},\frac{1}{2L}\}$.
B.2 Dynamic regret and proof of rest of the results in Section 3.1
@@
which can be further bounded by taking any $\bar{x}\in\mathcal{X}$ and
$\max_{t\in[T]}\max_{x\in\mathcal{X}}\|\nabla\ell_{t}(x)\|\leq\max_{t\in[T]}\|\nabla\ell_{t}(\bar{x})\|+LD.$
This constant $M$ will appear in a $\log\log$ term in the final result.
@@ B.2.1 Algorithm design
input initial vector $p^{1}=\tfrac{1}{N}\textbf{1}_{N}$ and $\eta_{i}=\eta_{\min}2^{i-1},i\in[N]$, exponentiated gradient learning rate $\delta>0$, initial point $x^{1}_{i}=x^{1},i\in[N]$.
@@ B.2.3 Algorithm analysis
$\textstyle\hat{\rho}(\{\hat{x}_{t}\}_{t=1}^{T})\leq{}$ $8\sqrt{(3+\log N){D^{2}}+2\mathcal{P}_{T}D}+8(3+\log N)LD^{2}+8LD\mathcal{P}_{T}$
$={}$ $\mathcal{O}(\sqrt{(1+\mathcal{P}_{T}+\hat{G}_{T})(1+\mathcal{P}_{T})}\log\log T)$
@@ Appendix C Proof of results in Section 4
$T[f(\bar{x}^{T})-f(x^{\star})]\leq{}$ $\textstyle\mathbb{E}\big[\sum_{t=1}^{T}[f(x^{t})-f(x^{\star})]\big]$
@@
C.2 Proof of Corollary 4.2
Combining Theorem 1 of [5] with the regret guarantees in Section˜2 and the online-to-batch argument from Corollary˜4.1 completes the proof.
C.3 Proof of Theorem 4.1
@@
since $\eta^{2}-\tfrac{\eta}{L}\leq 0$ for $\eta\leq\tfrac{1}{L}$. Next we bound $\sum_{j\in\{k:k\geq 2,\eta_{k}\geq\frac{1}{L}\}}(\eta^{2}_{j}-\tfrac{\eta_{j}}{L})\|\nabla\ell_{j}(x^{j})\|^{2}$ by taking $a_{0}=\|\nabla\ell_{1}(x^{1})\|^{2}$, $a_{j}=\|\nabla\ell_{j+1}(x^{j+1})\|^{2}$ and $g(u)=\tfrac{\alpha^{2}}{u}-\frac{\alpha}{L\sqrt{u}}$ in Lemma˜A.1 (note that $g$ is non-increasing for $u\leq(2L\alpha)^{2}$, which corresponds to $\eta\geq\frac{1}{2L}$ and is satisfied for $\eta_{k}\geq\frac{1}{L}$):
@@
C.4 Proof of Proposition 4.1
The construction modifies the standard OCO lower bound. Consider a 1D OCO problem with
$\ell_{t}(x)=\tfrac{1+\sigma_{t}}{2}g(x)+\tfrac{1-\sigma_{t}}{2}g(-x)$
where $g(x)=\bigg\{\begin{array}[]{ll}x+\tfrac{3}{2},&x\geq-1\\ \tfrac{1}{2}[x+2]_{+}^{2},&\text{else}\end{array}$ and $\sigma_{t}\in\{-1,1\}$ is a Rademacher random variable. It is easy to check that $g$ is $L=1$-smooth, convex, with $\inf g=0$. Moreover, for $\mathcal{X}=[-\delta,\delta]\subseteq[-1,1]$, we have $\ell_{t}(x)=\sigma_{t}x+\tfrac{3}{2}$ and therefore, the regret in terms of $\{\ell_{t}\}_{t=1}^{T}$ coincides with that of the 1D OCO lower bound since for $x^{t},x\in\mathcal{X}$, we have
$\rho_{T}(x)={}$ $\textstyle\sum_{t=1}^{T}\ell_{t}(x^{t})-\sum_{t=1}^{T}\ell_{t}(x)$
$={}$ $\textstyle\sum_{t=1}^{T}[\sigma_{t}x^{t}+\tfrac{3}{2}]-\sum_{t=1}^{T}[\sigma_{t}x+\tfrac{3}{2}]$
$={}$ $\textstyle\sum_{t=1}^{T}\sigma_{t}(x^{t}-x).$
Furthermore, we can upperbound $L^{\mathcal{X}}_{T}(x)$ by
$L^{\mathcal{X}}_{T}(x)=\sum_{t=1}^{T}\ell_{t}(x)-\min_{u\in\mathcal{X}}\ell_{t}(u)=\sum_{t=1}^{T}\sigma_{t}x-\min_{u\in[-\delta,\delta]}\sigma_{t}u\leq 2\delta T.$
Suppose an $\mathcal{O}(LD^{2}+D\sqrt{LL^{\mathcal{X}}_{T}(x^{\star})})$ upper bound holds, then with $D=2\delta$, it would imply
$\rho_{T}(x)\leq\mathcal{O}(4\delta^{2}+2\delta\sqrt{2\delta T})=\mathcal{O}(\delta^{2}+\delta^{3/2}\sqrt{T}).$
However, the lower bound of OCO implies $\rho_{T}(x)\geq\Omega(D\sqrt{T})=\Omega(\delta\sqrt{T})$, and taking $\delta=T^{-\beta},\beta\in(0,\tfrac{1}{3}]$ leads to an upper bound of $\mathcal{O}(T^{\frac{1-3\beta}{2}})$ and a lower bound of $\Omega(T^{\frac{1}{2}-\beta})$ that contradict each other. Dividing the upper and lower bounds completes the proof.
C.5 Proof of Proposition 4.2
@@
$\textstyle\sum_{t=1}^{T}\ell^{1/\gamma}_{t}(x)-\inf\ell^{1/\gamma}_{t}={}$ $\textstyle\sum_{t=1}^{T}\ell^{1/\gamma}_{t}(x)-\min_{u\in\mathcal{X}}\ell_{t}(u)$
$\leq{}$ $\textstyle\sum_{t=1}^{T}\ell^{1/\gamma}_{t}(x)-\min_{u\in\mathcal{X}}\ell_{t}(u)=L^{\mathcal{X}}_{T}(x),$
where the equality uses the fact that $\inf\ell^{1/\gamma}_{t}=\min_{u\in\mathcal{X}}\ell_{t}(u)$ and the last inequality uses $\ell(x)\leq\ell^{1/\gamma}(x)$ from (16). This completes the proof.
C.6 Hardness of achieving an $\mathcal{O}(D\sqrt{G^{\mathcal{X}}_{T}(x)})$ regret upper bound
We show that an $\mathcal{O}(LD^{2}+D\sqrt{G_{T}^{\mathcal{X}}(x)})$ upper bound is also generally not achievable.
Proposition C.1.
Let $\gamma=\Omega(1)$ be some constant. For any $\tau\in[0,\tfrac{1}{2})$, there exists a loss sequence $\{\ell_{t}\}_{t=1}^{T}$ satisfying A1, A2 such that for any online algorithm and $x\in\mathcal{X}$, $\tfrac{\rho_{T}(x)}{LD^{2}+D\sqrt{G_{T}^{\mathcal{X}}(x)}}=\Omega(T^{\tau})$.
Proof.
Consider the lower bound instance with $\mathcal{X}=[-\delta,\delta]$, $D=2\delta$, and $\ell_{t}(x)=g_{t}x$, where $g_{t}=\sigma_{t}$ is a Rademacher random variable. Each $\ell_{t}$ is linear and thus 0-smooth. Then
$\ell_{t}^{1/\gamma}(x)=\min_{z\in[-\delta,\delta]}\{g_{t}z+\tfrac{\gamma}{2}(z-x)^{2}\},$
and for $x\in\mathcal{X}$, we have
$\|\nabla\ell_{t}^{1/\gamma}(x)\|^{2}=\gamma^{2}|\Pi_{[-\delta,\delta]}[x-\tfrac{1}{\gamma}g_{t}]-x|^{2}\leq 4\gamma^{2}\delta^{2}.$
Hence, $G_{T}^{\mathcal{X}}(x)\leq 4\gamma^{2}\delta^{2}T$. The lower bound argument gives $\rho_{T}(x)\geq\Omega(\delta\sqrt{T})$, while the desired $\mathcal{O}(LD^{2}+D\sqrt{G_{T}^{\mathcal{X}}(x)})$ upper bound with $L=0$ yields
$\rho_{T}(x)=\mathcal{O}(\delta\gamma\sqrt{4\delta^{2}T})=\gamma\delta^{2}\sqrt{T}.$
Taking $\delta=T^{-\beta}$, we have lower bound of $\Omega(T^{\frac{1}{2}-\beta})$ and an upper bound of $\mathcal{O}(\gamma T^{\frac{1}{2}-2\beta})$ that contradict each other.